# Towards a Modern and Lightweight Rendering Engine for Dynamic Robotic Simulations

### v1

@@ Abstract
Interactive dynamic simulators are an accelerator for developing novel robotic control algorithms and complex systems involving humans and robots. In user training and synthetic data generation applications, a high-fidelity visualization of the simulation is essential. Visual fidelity is dependent on the quality of the computer graphics algorithms used to render the simulated scene. Furthermore, the rendering algorithms must be implemented on the graphics processing unit (GPU) to achieve real-time performance, requiring the use of a graphics application programming interface (API). This paper presents a performance-focused and lightweight rendering engine supporting the Vulkan graphics API. The engine is designed to modernize the legacy rendering pipeline of Asynchronous Multi-Body Framework (AMBF), a dynamic simulation framework used extensively for interactive robotics simulation development. This new rendering engine implements graphical features such as physically based rendering (PBR), anti-aliasing, and ray-traced shadows, significantly improving the image quality of AMBF. Computational experiments show that the engine can render a simulated scene with over seven million triangles while maintaining GPU computation times within two milliseconds.
@@ I Introduction
Interactive dynamic simulators are an accelerator for developing novel robotic control algorithms and complex systems involving humans and robots. Specifically, dynamic simulators are cost-effective solutions to train human operators and machine learning algorithms in a wide range of domains. Effective simulation frameworks for robotics are required to produce realistic visualizations of the workspace while maintaining real-time performance. In previous work, the Asynchronous Multi-body framework (AMBF) [1] was introduced as a general real-time dynamic simulator for robots and free bodies. This framework was developed by integrating several external tools, such as an extended version of CHAI-3D, BULLET-Physics, and OpenGL. With its flexible description format, AMBF has enabled the development of complex interactive simulations, such as an environment for learning robot-assisted surgical suturing [2] and a volumetric simulation for skull base surgery [3].
@@
Currently, AMBF rendering and computing capabilities are limited by the legacy OpenGL graphics application programming interface (API) at the core of its rendering pipeline. AMBF relies on OpenGL version 2.1, which was released in 2006, to perform graphics operations. This is significant because it severely limits AMBF’s access to new developments in GPU hardware and real-time rendering that improve performance and visual fidelity [4, 5]. A particular limiting factor of legacy OpenGL is the lack of support for general-purpose computing on the GPU, a groundbreaking development that can greatly accelerate computationally intensive processes, such as soft-body physics simulation, ray tracing, and deep learning [6].
@@
To address these limitations, this work presents a new standalone rendering engine supporting the Vulkan graphics API to modernize AMBF, which may be applicable to other simulators limited by legacy rendering engines. We chose Vulkan, OpenGL’s successor, due to its cross-platform support and low-level design, allowing for significant performance improvements. Given the substantial architectural differences between legacy OpenGL and Vulkan, it was decided to develop the renderer as a standalone program and, therefore, integrate both processes via a plugin in AMBF. As such, support for interface devices, physics computations, and plugin architecture is maintained in AMBF, while the proposed engine renders the geometry of the scene (see Figure 1). Overall, the full system enables access to all AMBF features while significantly improving rendering quality.
@@
While developing this newer engine, efforts were concentrated on improving the computational performance of the rendering pipeline compared to the previous version of AMBF. Additionally, several additional features, such as support for physically-based rendering (PBR) materials, anti-aliasing, and ray-traced shadows, were co-developed to improve the visual fidelity of the engine. Overall, this paper summarizes some of the main design principles and challenges faced while developing this system, as well as opportunities for using a modern graphics API in developing advanced dynamic simulations. In summary, our key contributions are:
@@
    A standalone and open-source rendering engine supporting the Vulkan graphics API.
@@
    Integration between the AMBF dynamic simulation engine and the standalone rendering engine, as an illustration of how this approach could be used with other simulators.
Figure 1: Overview of the proposed dynamic simulation framework. User input is communicated to AMBF through interface devices. AMBF Simulator runs physics simulations on the scene and records scene updates in shared memory. AMBF-Vulkan reads the scene updates from shared memory and renders the result with high visual quality.
@@ II Related Work
For AMBF, our previous work [9] presented a plugin that enabled rendering AMBF scenes using Blender’s built-in rendering engines, EEVEE and Cycles[10], and using ROS [11] as a communication layer. This plugin permitted AMBF to use all rendering features provided by Blender; however, it could not render scenes at interactive rates due to communication latency and the processing overhead of Blender’s interface. Compared to our previous work, the proposed simulation framework presents several optimizations that simultaneously improve visual fidelity and rendering speed. The framework’s overall performance can be attributed to the use of a modern graphics API and a lightweight communication layer between AMBF and the rendering engine.
The Vulkan API caught our attention due to its great performance on ray tracing [12, 13], and complex shading advancements [14] that could facilitate advanced visual effects such as light reflection and detailed material textures. Furthermore, the Vulkan API is cross-platform [15, 16] and facilitates the use of ray tracing acceleration units[17]. Vulkan has been increasingly implemented in recent academic research [18]. Taking advantage of the low-level API and integration compatibility, Vulkan has been deployed into various domains: physics particle simulation [19, 20], efficient data visualization [21], deformable body simulation [22], and surgical simulations [23, 24]. Given all the advantages presented, Vulkan is an excellent option for modern rendering engines.
Figure 2: Overview of Vulkan’s render loop from the scene update to the presentation of the image to the screen. The scene geometry is rendered in the Main Pass, and its resulting image is used as a sampled texture by the Post-Processing Pass. The ImGui pass renders the GUI overtop the image and the result is presented to the screen.
@@ III-A Overview of graphics APIs and OpenGL
Modern OpenGL (versions 3.0 to 4.6) progressively shifted towards a more programmable graphics pipeline, adding support for features like programmable shaders, compute shaders, and multi-threaded command buffer submission, among others. In 2017, active development of OpenGL was ceased in favor of the newer Vulkan API, making OpenGL 4.6 the final version. This is relevant as dynamic simulators relying on OpenGL will never receive support for new GPU features, such as hardware-accelerated ray tracing. Moreover, GPU manufacturers and driver developers may decide to stop supporting the OpenGL specification, as seen by Apple deprecating OpenGL on all of their platforms in 2018. Overall, it was decided to migrate AMBF directly to Vulkan instead of a newer OpenGL version to ensure long-term support for the simulator in modern systems.
@@ IV Methodology
The complete dynamic simulation framework is composed of three different components (see Fig. 1). Firstly, we have the standalone rendering engine built for the Vulkan API, which we will refer to from this point on as AMBF-Vulkan. By itself, the AMBF-Vulkan engine only renders the loaded objects and provides the option to move the virtual cameras and adjust visual parameters via a GUI. Secondly, we have AMBF running in headless mode to perform physics computations and communicate with user input devices. Thirdly, we have the communication plugin that enables AMBF to update the pose of all the objects inside the AMBF-Vulkan engine via shared memory between both processes. Overall, these three components work together to enable the user to access the full set of features from the dynamic simulator AMBF, while also obtaining high-quality renderings from the Vulkan rendering engine. The rest of the methodology is divided as follows. Subsection IV-A describes the software design of the renderer AMBF-Vulkan. Subsection IV-B presents the graphical features that were implemented in AMBF-Vulkan to improve visual fidelity. Lastly, subsection IV-C describes the communication plugin between AMBF and AMBF-Vulkan.
IV-A AMBF-Vulkan engine rendering architecture
@@ IV-A1 Render Loop Overview
Figure 2 illustrates a high-level overview of the AMBF-Vulkan engine’s rendering loop. The data flow depicted is as follows: (1) The scene is updated to reflect the changes made from device input and model transformation updates (see IV-C), (2) mesh draw commands are sorted and submitted for rendering, (3) post-processing operations are applied to the main pass’s output, (4) the Dear ImGui [31] library renders an immediate-mode Graphical User Interface (GUI) onto the post-processed image, and (5) the final frame is presented to the screen.
@@
Key to the performance of any rendering engine is its implementation for storing scene assets and submitting GPU draw commands. AMBF-Vulkan opts to store model resource handles in global type-specific containers. The scene is merely a nodal hierarchy, with each node pointing to the resources it uses. This way, we can create a collection of draw command requests containing the necessary resource information and sort the list to optimize draw performance. We accomplish this by sorting by material type and vertex buffer index, which minimizes the amount of descriptor (a modern graphics API term to describe bound shader resources, e.g., textures or uniform buffers) bind commands and vertex buffer cache misses.
@@
With the draw commands sorted, we can submit the Main Pass command buffer to the GPU. The vertex shader of the main pass is responsible for transforming the vertices into view space and passing shading data to the fragment shader. AMBF-Vulkan uses a performance-improving technique called vertex pulling: rather than additionally binding vertex buffers for each draw, we store the vertex information in a buffer and index into it, like an array. The fragment shader is responsible for lighting calculations discussed in detail in IV-B. Additionally, the fragment shader converts color data from sRGB gamma space to linear space to perform lighting calculations. Afterward, we apply Reinhard tone mapping to map color values greater than 1.0 to a value between 0.0 and 1.0. Otherwise, they will be clamped to 1.0 automatically. Finally, we invert the aforementioned linear space conversion to bring the colors back into sRGB gamma space, an operation known as gamma correction.
@@
The image result of the Main Pass is then converted for use as a texture image in the following pass, the Post-Processing Pass. This render pass is typically used to perform image processing techniques on a rendered image. In AMBF-Vulkan, we currently use the Post-Processing Pass to perform a screen-space anti-aliasing algorithm discussed in IV-B2. The vertex shader creates a screen-sized quad and passes its UV indices to the fragment shader. The fragment shader samples the Main Pass resulting texture and performs the anti-aliasing algorithm.
@@
Lastly, the result of the Post-Processing Pass is converted into an image format matching that of the display device. The blitted image is then used to render a GUI over the scene using the Dear ImGui library. Rather than a retained mode GUI, which can add a lot of unnecessary complexity due to state management, we opted for an immediate mode GUI. While this choice adds unnecessary drawing of unchanged GUI state, the cost is negligible for our purposes, as seen in Figure 7. In return, we get the flexibility and simplicity of the Dear ImGui library for our debugging and scene customization needs. Currently, the GUI displays a window listing basic debugging and profiling information, e.g., frame time and camera coordinates. With the ImGui Pass finished, the final image is ready to be presented to the screen.
@@ IV-A2 System Improvements
Rather than an uncapped frame rate, the system synchronizes with the display device, limiting the frame rate to the refresh rate of the display device. This prevents unnecessary computation while allowing the rate cap to be display-dependent, which can be beneficial if the simulation is displayed via a virtual reality headset. AMBF-Vulkan uses a double-buffered swapchain, i.e., while one frame is rendered by the GPU, the other is prepared for rendering by the CPU. As shown in Figure 2, render attachments, the resources that the render pass writes to, are recycled from the previous frame. In actuality, this means two frames prior due to the double buffer.
@@
IV-B AMBF-Vulkan graphical features
A central motivation for building a modernized rendering engine for AMBF was to surpass the graphical quality restrictions of our previous implementation. In the context of surgical simulations, realistic visualizations play an important role in both operator training and machine learning. The following section describes the graphical improvements that were implemented in AMBF-Vulkan.
@@ IV-B1 Physically Based Rendering
Physically Based Rendering (PBR) is a lighting model approach in Computer Graphics that enables photo-realism through approximation of real-world physics[32]. This is especially useful in a simulation environment where real-world measurements can be used to create a more accurate representation. Figures 3a, 3b and 3c demonstrate how we can use normal incidence Fresnel reflectance measurements to shade metallic materials more accurately. Figures 3d and 3e demonstrate different roughness values on a non-metallic material. AMBF-Vulkan uses a Cook-Torrance specular reflectance model variation that is common in real-time rendering applications [33]:
@@ IV-B2 Anti-Aliasing
AMBF-Vulkan employs two techniques to reduce visual artifacts caused by under-sampling the scene (see Figure 4). These techniques differ in the types of aliasing they target, as well as the stage of rendering they are implemented in, and we combine them to improve image quality further.
Multisample Anti-Aliasing (MSAA) allows for additional sampling per pixel with little performance cost. Rather than shading each additional sample, as with Supersampling Anti-Aliasing, additional samples are only shaded if the triangle they intersect is not oversampled within the pixel. Before beginning the main render pass, AMBF-Vulkan submits a multisample draw and depth image, along with an additional resolve draw and depth image, and the GPU will automatically handle the rest.
@@ IV-B3 Ray-traced shadows
GPU ray tracing support is a clear example of the benefits of moving to a modern graphics API such as Vulkan. Similar to PBR, ray tracing approximates real-world light transport physics, resulting in more photorealistic images. Due to its high computational expense, ray tracing in real-time applications was limited until recent advancements in hardware acceleration and denoising techniques. Using Vulkan’s ray-tracing extensions, we implemented dynamic ray-traced hard shadows as an alternative to shadow mapping techniques (See Figure 5).
Ray-traced shadow requires the implementation of an acceleration structure build system. In short, acceleration structures enable the computation of ray intersections with scene geometry in logarithmic time complexity and are essential for real-time ray tracing. Once the acceleration structures are set in place, we can use the GLSL_EXT_ray_query extension from the fragment shader to ray-trace the scene and get important information, such as whether or not a pixel is in shadow. To determine if a pixel is in shadow, a ray is traced from the fragment’s world space coordinate to the light’s world space coordinate. If the ray hits any opaque geometry during the tracing, then the fragment is in shadow and does not receive direct light.
To build the acceleration structures in our engine the following steps are performed: (1) on scene initialization, the geometry of each unique mesh is built into a corresponding GPU-side buffer called a Bottom Level Acceleration Structure (BLAS), (2) each BLAS is subsequently compacted for faster traversal and a smaller memory footprint, (3) an Instance object is built for each mesh instance, recording its initial transformation matrix, (4) the transforms are updated on each frame’s scene update, and a single Top Level Acceleration Structure (TLAS) is built, containing each Instance. Because the application does not yet support geometry deformation, updating or rebuilding the BLASs is unnecessary. Furthermore, because TLAS updates deteriorate scene traversal quality, we choose to rebuild the TLAS every frame for a net increase in performance. Additionally, we offload all acceleration structure build tasks to an async compute queue, virtually removing the performance cost.
@@
IV-C Plugin communication between AMBF and AMBF-Vulkan
To obtain a complete dynamics simulator, the rendering engine AMBF-Vulkan requires a physics engine that can update the poses of all the bodies in the simulation. In this work, we choose the physics engine inside AMBF, but the software design is flexible enough that other engines could be used. To enable communication between AMBF and the Vulkan engine, we designed the AMBF-Vulkan Interchange (AVI) plugin, which enables AMBF to write the pose of all the bodies at every physics iteration into shared memory. AMBF-Vulkan then uses the poses stored in shared memory before rendering the scene (see Fig. 6).
Figure 6: Overview of the engine interchange system. The scene is exported from Blender to a glTF file and, using the AMBF Blender add-on, an ADF file. The ADF file is loaded in AMBF Simulator to communicate with interface devices and simulate physics. The AVI plugin records the model transform matrices into shared memory every frame. AMBF Vulkan loads the glTF file and reads from shared memory every frame to update each model’s pose.
AVI uses the boost::interprocess library for a fast, lightweight, and operating-system-agnostic solution for interprocess communication. We simply use AVI to record AMBF’s copy of each mesh node’s world-space transformation matrix in a shared memory region, and AMBF-Vulkan updates its own copy of the mesh node’s matrix. This one-way data transfer approach avoids many of the pitfalls of concurrency. Additionally, the library’s managed_shared_memory class will internally handle the ownership swapping of the shared memory region using a recursive mutex. Lastly, we chose to simply overwrite each matrix every frame and read each matrix every frame, regardless of whether or not the data has changed. We can afford to use this inefficient but error-resistant implementation because of the relatively low amount of meshes in a typical AMBF simulation environment.

### v2

# FIRE-3DV: Framework-Independent Rendering Engine for 3D Graphics using Vulkan

@@ Abstract
Interactive dynamic simulators are an accelerator for developing novel robotic control algorithms and complex systems involving humans and robots. In user training and synthetic data generation applications, high-fidelity visualizations from the simulation are essential. Yet, robotic simulators often limit their rendering algorithms to preserve real-time interaction with the simulation. Advancements in Graphics Processing Units (GPU) enable improved visualization without compromising performance. However, these advancements cannot be fully leveraged in simulation frameworks that use legacy graphics application programming interfaces (API) to interface with the GPU. This paper presents a performance-focused and lightweight rendering engine supporting the modern Vulkan graphics API that can be easily integrated with other simulation frameworks to enhance visualizations. To illustrate the proposed method, our engine is used to modernize the legacy rendering pipeline of the Asynchronous Multi-Body Framework (AMBF), a dynamic simulation framework used extensively for interactive robotics simulation development. This new rendering engine implements graphical features such as physically based rendering (PBR), anti-aliasing, and ray-traced shadows, significantly improving the image fidelity of AMBF. Computational experiments show that the engine can render a simulated scene with over seven million triangles while maintaining GPU computation times within two milliseconds.
@@ I Introduction
Dynamic robotic simulators facilitate the development and testing of novel robotic control algorithms and complex systems involving humans and robots. In particular, interactive dynamic simulators have been used extensively as cost-effective and safe solutions to train human operators and machine learning algorithms in a wide range of domains [1]. Simulation frameworks targeted for training applications require realistic visualizations of the workspace while maintaining real-time performance. In previous work, the Asynchronous Multi-Body Framework (AMBF) [2] was introduced as a general real-time dynamic simulator for robots and free bodies. This framework was developed by integrating several external tools, such as an extended version of CHAI-3D, BULLET-Physics, and OpenGL. With its flexible description format, AMBF has enabled the development of complex interactive simulations, such as an environment for learning robot-assisted surgical suturing [3] and a volumetric simulation for skull base surgery [4].
@@
Currently, AMBF rendering and computing capabilities are limited by the legacy OpenGL graphics application programming interface (API) at the core of its rendering pipeline. AMBF relies on OpenGL version 2.1, released in 2006, to perform graphics operations. This severely limits AMBF’s access to new developments in GPU hardware and real-time rendering that improve performance and visual fidelity [5, 6]. A particular limiting factor of legacy OpenGL is the lack of support for general-purpose computing on the GPU, a groundbreaking development that can greatly accelerate computationally intensive processes, such as soft-body physics simulation, ray-tracing, and deep learning [7]. Other simulation engines dependent on OpenGL face similar issues and will eventually need to pivot to a modern graphics API to stay compatible and performant on modern hardware.
@@
This work presents a framework-independent rendering engine for 3D graphics using Vulkan (FIRE-3DV) to modernize AMBF and other simulators limited by legacy rendering engines. We chose Vulkan, OpenGL’s successor, due to its cross-platform support and low-level design, allowing for significant performance improvements. Given the substantial architectural differences between legacy OpenGL and Vulkan, it was decided to develop the renderer as a standalone program and, therefore, integrate both processes via a plugin in the simulation engine. As such, support for interface devices, physics computations, and plugin architecture is maintained in the simulation engine, while FIRE-3DV only takes care of the scene visualization (see Fig. 1). Overall, the full system retains access to the simulation engine’s features while significantly improving rendering quality.
@@
While developing this rendering engine, efforts were concentrated on improving the computational performance of the rendering pipeline compared to AMBF. Additionally, several additional features, such as support for physically-based rendering (PBR) materials, anti-aliasing, and ray-traced shadows, were co-developed to improve the visual fidelity of the engine. Overall, this paper summarizes some of the main design principles and challenges faced while developing this system, as well as opportunities for using a modern graphics API in developing advanced dynamic simulations. In summary, our key contributions are:
@@
    FIRE-3DV, a standalone and open-source rendering engine supporting the Vulkan graphics API.
@@
    Integration between the AMBF dynamic simulation engine and FIRE-3DV, as an illustration of how this approach could be used with other simulators.
Figure 1: Overview of the proposed dynamic simulation framework. The simulation engine handles user input through interface devices and runs physics simulations on the scene, recording scene updates in shared memory. FIRE-3DV reads the scene updates from shared memory and renders the result with high visual fidelity.
@@ II Related Work
For AMBF, our previous work [10] presented a plugin that enabled rendering AMBF scenes using Blender’s built-in rendering engines, EEVEE and Cycles[11], and used the Robotic Operating System (ROS) [12] as a communication layer. This plugin permitted AMBF to use all rendering features provided by Blender; however, it could not render scenes at interactive rates due to communication latency and the processing overhead of Blender’s interface. Compared to our previous work, the proposed rendering engine presents several optimizations that simultaneously improve visual fidelity and rendering speed. The systems’s overall performance can be attributed to the use of a modern graphics API and a lightweight communication layer between AMBF and the proposed rendering engine.
The Vulkan API caught our attention due to its great performance on ray-tracing [13, 14], and complex shading advancements [15] that could facilitate advanced visual effects such as light reflection and detailed material textures. Furthermore, the Vulkan API is cross-platform [16, 17] and facilitates the use of ray-tracing acceleration units[18]. Vulkan has been increasingly implemented in recent academic research [19]. Taking advantage of the low-level API and integration compatibility, Vulkan has been deployed into various domains: physics particle simulation [20, 21], efficient data visualization [22], deformable body simulation [23], and surgical simulations [24, 25]. Given all the advantages presented, Vulkan is an excellent option for modern rendering engines.
Figure 2: Overview of FIRE-3DV’s render loop from the scene update to the presentation of the image to the screen. The scene geometry is rendered in the Main Pass, and its resulting image is used as a sampled texture by the Post-Processing Pass. The ImGui pass renders the GUI overtop the image and the result is presented to the screen.
@@ III-A Overview of graphics APIs and OpenGL
Modern OpenGL (versions 3.0 to 4.6) progressively shifted towards a more programmable graphics pipeline, adding support for features like programmable shaders, compute shaders, and multi-threaded command buffer submission, among others. In 2017, active development of OpenGL was ceased in favor of the newer Vulkan API, making OpenGL 4.6 the final version. This is relevant as dynamic simulators relying on OpenGL will never receive support for new GPU features, such as hardware-accelerated ray-tracing. Moreover, GPU manufacturers and driver developers may decide to stop supporting the OpenGL specification, as seen by Apple deprecating OpenGL on all of their platforms in 2018. Overall, it was decided to support Vulkan instead of a newer OpenGL version to ensure long-term support for the rendering engine on modern hardware.
@@ IV Methodology
The complete dynamic simulation framework is composed of three different components (see Fig. 1). Firstly, we have the standalone rendering engine Framework-Independent Rendering Engine for 3D graphics using Vulkan (FIRE-3DV). By itself, FIRE-3DV only renders the loaded objects and provides the option to move the virtual cameras and adjust visual parameters via a GUI. Secondly, we have the simulation engine running in headless mode to perform physics computations and communicate with user input devices. Thirdly, we have a memory region shared between both processes that is used to record scene updates given by the simulation engine to update the scene in FIRE-3DV. Overall, these three components work together to enable the user to access the full set of features from the dynamic simulator, while also obtaining high-quality renderings from FIRE-3DV. The rest of the methodology is divided as follows. Subsection IV-A describes the software design of the renderer FIRE-3DV. Subsection IV-B presents the graphical features that were implemented in FIRE-3DV to improve visual fidelity. Lastly, subsection IV-C describes the interprocess communication plugin for AMBF and FIRE-3DV as example for how other simulators could integrate with FIRE-3DV.
IV-A FIRE-3DV rendering engine architecture
@@ IV-A1 Render Loop Overview
Figure 2 illustrates a high-level overview of the rendering loop implemented in FIRE-3DV. The data flow depicted is as follows: (1) The scene is updated to reflect the changes made from device input and model transformation updates (see IV-C), (2) mesh draw commands are sorted and submitted for rendering, (3) post-processing operations are applied to the main pass’ output, (4) the Dear ImGui [32] library renders an immediate-mode Graphical User Interface (GUI) onto the post-processed image, and (5) the final frame is presented to the screen.
@@
Key to the performance of our rendering engine is its implementation for storing scene assets and submitting GPU draw commands. FIRE-3DV opts to store model resource handles in global type-specific containers. The scene data is organized by a nodal hierarchy, with each node pointing to the resources it uses. This way, we can create a collection of draw command requests containing the necessary resource information and sort the list to optimize draw performance. We accomplish this by sorting by material type and vertex buffer index, which minimizes the amount of descriptor (a modern graphics API term to describe bound shader resources, e.g., textures or uniform buffers) bind commands and vertex buffer cache misses.
@@
With the draw commands sorted, we can submit the Main Pass command buffer to the GPU. The vertex shader of the main pass is responsible for transforming the vertices into view space and passing shading data to the fragment shader. FIRE-3DV uses a performance-improving technique called vertex pulling: rather than additionally binding vertex buffers for each draw, we store the vertex information in a buffer and index into it, like an array. The fragment shader is responsible for lighting calculations discussed in detail in IV-B. Additionally, the fragment shader converts color data from sRGB gamma space to linear space to perform lighting calculations. Afterward, we apply Reinhard tone mapping to map color values greater than 1.0 to a value between 0.0 and 1.0. Otherwise, they will be clamped to 1.0 automatically. Finally, we invert the aforementioned linear space conversion to bring the colors back into sRGB gamma space, an operation known as gamma correction.
@@
The image result of the Main Pass is then converted for use as a texture image in the following pass, the Post-Processing Pass. This render pass is typically used to perform image processing techniques on a rendered image. In FIRE-3DV, we currently use the Post-Processing Pass to perform a screen-space anti-aliasing algorithm discussed in IV-B2. The vertex shader creates a screen-sized quad and passes its UV indices to the fragment shader. The fragment shader samples the Main Pass resulting texture and performs the anti-aliasing algorithm.
@@
Lastly, the result of the Post-Processing Pass is converted into an image format matching that of the display device. The blitted image is then used to render a GUI over the scene using the Dear ImGui library. Rather than a retained-mode GUI, which can add a lot of unnecessary complexity due to state management, we opted for an immediate-mode GUI. While this choice adds unnecessary drawing of unchanged GUI state, the cost is negligible for our purposes, as seen in Figure 7. In return, we get the flexibility and simplicity of the Dear ImGui library for our debugging and scene customization needs. Currently, the GUI displays a window listing basic debugging and profiling information, e.g., frame time and camera coordinates. With the ImGui Pass finished, the final image is ready to be presented to the screen.
@@ IV-A2 System Improvements
Rather than an uncapped frame rate, the system synchronizes with the display device, limiting the frame rate to the refresh rate of the display device. This prevents unnecessary computation while allowing the rate cap to be display-dependent, which can be beneficial if the simulation is displayed via a virtual reality headset. FIRE-3DV uses a double-buffered swapchain, i.e., while one frame is rendered by the GPU, the other is prepared for rendering by the CPU. As shown in Figure 2, render attachments, the resources that the render pass writes to, are recycled from the previous frame. In actuality, this means two frames prior due to the double buffer.
@@
IV-B Graphical features of FIRE-3DV
A central motivation for building a modernized rendering engine for dynamic simulation was to surpass the graphical quality restrictions of our previous implementation, AMBF. In the context of surgical simulations, realistic visualizations play an important role in both operator training and machine learning. The following section describes the graphical improvements that were implemented in FIRE-3DV.
@@ IV-B1 Physically Based Rendering
Physically Based Rendering (PBR) is a lighting model approach in Computer Graphics that enables photo-realism through approximation of real-world physics[33]. This is especially useful in a simulation environment where real-world measurements can be used to create a more accurate representation. Figures 3a, 3b and 3c demonstrate how we can use normal incidence Fresnel reflectance measurements to shade metallic materials more accurately. Figures 3d and 3e demonstrate different roughness values on a non-metallic material. FIRE-3DV uses a Cook-Torrance specular reflectance model variation that is common in real-time rendering applications [34]:
@@ IV-B2 Anti-Aliasing
FIRE-3DV employs two techniques to reduce visual artifacts caused by under-sampling the scene (see Figure 4). These techniques differ in the types of aliasing they target, as well as the stage of rendering they are implemented in, and we combine them to improve image quality further.
Multisample Anti-Aliasing (MSAA) allows for additional sampling per pixel with little performance cost. Rather than shading each additional sample, as with Supersampling Anti-Aliasing, additional samples are only shaded if the triangle they intersect is not oversampled within the pixel. Before beginning the main render pass, FIRE-3DV submits a multisample draw and depth image, along with an additional resolve draw and depth image, and the GPU will automatically handle the rest.
@@ IV-B3 Ray-traced shadows
GPU ray-tracing support is a clear example of the benefits of moving to a modern graphics API such as Vulkan. Similar to PBR, ray-tracing approximates real-world light transport physics, resulting in more photorealistic images. Due to its high computational expense, ray-tracing in real-time applications was limited until recent advancements in hardware acceleration and denoising techniques. Using Vulkan’s ray-tracing extensions, we implemented dynamic ray-traced hard shadows as an alternative to shadow mapping techniques (See Figure 5).
Algorithms designed to utilize ray-tracing acceleration hardware require the implementation of an acceleration structure build system. In short, acceleration structures enable the computation of ray intersections with scene geometry in logarithmic time complexity and are essential for real-time ray-tracing. Once the acceleration structures are set in place, we can use the GLSL_EXT_ray_query extension from the fragment shader to ray-trace the scene and get important information, such as whether or not a pixel is in shadow. To determine if a pixel is in shadow, a ray is traced from the fragment’s world space coordinate to the light’s world space coordinate. If the ray hits any opaque geometry during the tracing, then the fragment is in shadow and does not receive direct light.
To build the acceleration structures in our engine, the following steps are performed: (1) on scene initialization, the geometry of each unique mesh is built into a corresponding GPU-side buffer called a Bottom Level Acceleration Structure (BLAS), (2) each BLAS is subsequently compacted for faster traversal and a smaller memory footprint, (3) an Instance object is built for each mesh instance, recording its initial transformation matrix, (4) the transforms are updated on each frame’s scene update, and a single Top Level Acceleration Structure (TLAS) is built, containing each Instance. Because the application does not yet support geometry deformation, updating or rebuilding the BLASs is unnecessary. Furthermore, because TLAS updates deteriorate scene traversal quality, we choose to rebuild the TLAS every frame for a net increase in performance. Additionally, we offload all acceleration structure build tasks to an async compute queue, virtually removing the performance cost.
@@
IV-C Plugin communication between AMBF and FIRE-3DV
To obtain a complete dynamics simulator, the rendering engine FIRE-3DV requires a physics engine that can update the poses of all the bodies in the simulation. In this work, we choose the physics engine inside AMBF, but the software design is flexible enough that other engines could be used. To enable communication between AMBF and FIRE-3DV, we designed the AMBF-FIRE-3DV Interchange (AFI) plugin, which enables AMBF to write the pose of all the bodies at every physics iteration into shared memory. FIRE-3DV then uses the poses in shared memory before rendering the scene (see Fig. 6).
Figure 6: Overview of the AMBF-FIRE-3DV interchange system. The scene is exported from Blender and then imported into AMBF and FIRE-3DV. The AFI plugin records the transform matrices from all the objects inside AMBF and writes them into shared memory at every physics update. FIRE-3DV reads from shared memory to update each model’s pose before rendering.
AFI uses the boost::interprocess library for a fast, lightweight, and operating-system-agnostic solution for interprocess communication. We simply use AFI to record AMBF’s copy of each mesh node’s world-space transformation matrix in a shared memory region, and FIRE-3DV updates its own copy of the mesh node’s matrix. This one-way data transfer approach avoids many of the pitfalls of concurrency. Additionally, the library’s managed_shared_memory class will internally handle the ownership swapping of the shared memory region using a recursive mutex. Lastly, we chose to simply overwrite each matrix every frame and read each matrix every frame, regardless of whether or not the data has changed. We can afford to use this inefficient but error-resistant implementation because of the relatively low amount of meshes in a typical AMBF simulation environment.
IV-D Integration with other dynamic simulation engines
Similarly to AMBF, other 3D dynamic simulators can intercommunicate with FIRE-3DV to achieve higher-quality visual results with little modification to the simulation engine. Using a plugin architecture or inserting a small amount of code into the render loop, simulation engines must write each models transformation matrix into shared memory every frame, allowing FIRE-3DV to update the poses. Optionally, the simulation engine can write a change of basis matrix into shared memory to account for differences in bases between the simulation engine and FIRE-3DV. If the simulation engines uses a different file format than glTF, the scene can be exported to glTF using a tool like Blender, enabling FIRE-3DV to load the scene.